# The Redshift Evolution of the $M_{\bullet}-M_{\star}$ Relation for JWST’s Supermassive Black Holes at $z>4$

### 08 Jan 2024 v1

@@ 1 Introduction
Other works have studied the redshift evolution of scaling relations using numerical simulations (e.g., Robertson et al. 2006; Di Matteo et al. 2008; Sijacki et al. 2015), semi-analytic models (e.g., Malbon et al. 2007; Kisaka & Kojima 2010), observations (e.g., Peng et al. 2006; Decarli et al. 2010; Merloni et al. 2010; Trakhtenbrot & Netzer 2010; Bennert et al. 2011), or combinations of those (e.g., Booth & Schaye 2011), especially at $z\lesssim 2$. More recently, Caplar et al. (2018) developed a phenomenological model, based on observations, and found that the ratio $M_{\bullet}/M_{\star}$ scales as $\propto(1+z)^{1.5}$ at $z<2$, in agreement with Wyithe & Loeb (2003).
@@ 2 Model Principles
The basis of our model is that the black hole mass is the primary parameter that controls high-$z$ star formation quenching. This premise is supported by recent analyses of JWST/CEERS data with cosmological simulations (Illustris TNG and EAGLE), showing that high-$z$ galaxy quenching is primarily regulated by the mass of the SMBH (Piotrowska et al., 2022; Bluck et al., 2023). Previous cosmological simulations already showed that star formation quenching should scale with energy input from the central SMBH over the entire lifetime of the galaxy, which is proportional to the black hole mass (Terrazas et al., 2020; Bluck et al., 2023). Active SMBHs in $z>4$ galaxies discovered by JWST are effective at quenching star formation for at least two reasons.
First, high-$z$ galaxies are physically small; the ionized bubble generated by active SMBHs likely extends to the entire galaxy. Typical physical sizes of galaxies detected by JWST at $z=7-9$ are characterized by effective radii of $80\,\mathrm{pc}<r_{e}<300\,\mathrm{pc}$, with a mean value of $r_{e}\sim 150$ pc (Baggen et al., 2023). This typical physical size has to be compared with the radius of the ionization bubble created by a SMBH accreting at its Eddington rate. For a $\sim 10^{7}\,{\rm M_{\odot}}$ SMBH, as typically found in these overmassive systems (Pacucci et al., 2023), this radius can extend as much as $\sim 700$ kpc (see, e.g., Cen & Haiman 2000; Madau & Rees 2000; White et al. 2003). Regions of high-density gas presumably present in the high-$z$ galaxies could effectively shield the radiation from the SMBH and still allow localized star formation; this would, however, be ineffective in generating large-scale star formation. Hence, SMBHs were effective in quenching star formation because they had a global impact on the entire host. Note that previous studies (e.g., Chen et al. 2020) already highlighted the importance of the radius of star-forming galaxies in determining the growth of their SMBHs.
@@
Second, in the high-$z$ Universe, the growth time of black holes is comparable to the age of the Universe. Hence, central SMBHs at a given redshift $z$ are active for a time comparable to the Hubble time $t_{age}(z)$, if $z>4$. Figure 1 shows, for the high-$z$ overmassive systems detected thus far by JWST, a comparison between the age of the Universe (at detection) and growth time. This latter time is calculated assuming a continuous growth from a seeding redshift of $z=25$ (see, e.g., Barkana & Loeb 2001) at the Eddington rate, assuming a light seed of $100\,{\rm M_{\odot}}$ (left panel) or a heavy seed of $10^{4}\,{\rm M_{\odot}}$ (right panel). At higher redshift, the age of the Universe is comparable to, or even shorter than, the growth time (for those two specific seeding scenarios); hence, the “activity duty cycle” for those specific SMBHs has to be close to unity. Because their duty cycle is close to unity, these SMBHs are constantly injecting energy into the primeval galaxy and heating the cold gas necessary to produce stars. Once the age of the Universe is $\gtrsim 1$ Gyr, the SMBH is active only for a fraction of the Hubble time and stars can then form from cold molecular gas, which becomes widely available in the galaxy.
This hypothesis is further confirmed by a recent analysis of $124.5<z<$ 12 galaxies in the JWST/CEERS survey (Cole et al., 2023), showing a higher variability of star formation activity at high redshift. Stars form primarily in short periods of starburst activity, with star-forming duty cycles of only $20\%$ at $z\sim 9$, and $40\%$ at $z\sim 5$. The study also suggests a smoother star formation activity at $z<4.5$, when the age of the Universe is $>1\,\rm Gyr$ and the quasar duty cycle drops significantly below unity.
@@ 3.1 A Condition for Star Formation Quenching
We assume that star formation is quenched once the SMBH injects sufficient energy into the system to raise the temperature above the virial one. Recently, Gelli et al. (2023) used a similar formalism to argue that supernova feedback fails to quench star formation in high-$z$ galaxies. This finding supports our model principles detailed in Sec. 2.
@@
We adopt $\,{f_{\rm Edd}}=1$ for our calculations because overmassive systems discovered by JWST are estimated to be accreting at rates $50.1<\,{f_{\rm Edd}}<$ 5 (Harikane et al., 2023; Maiolino et al., 2023a). Furthermore, we use the median cooling function for the gas metallicity range $0.30.1<Z/Z_{\odot}<$ 0.3 and the mean gas number density ($\sim 0.5\,\mathrm{cm^{-3}}$) calculated for simulated galaxies in the redshift range $105<z<$ 10 by Robinson et al. (2022). The metallicity range used is justified by a recent study with JWST of $z\sim 6$ galaxies with masses $\sim 10^{10}\,{\rm M_{\odot}}$, showing typical values $12+\log(\mathrm{O/H})\sim 8.2$, which is $\approx 25\%$ of the solar value (Nakajima et al., 2023). Finally, we assume the values of the cosmological parameters from Planck Collaboration et al. (2020) and the halo mass to stellar mass ratio from Behroozi et al. (2019).
@@
Figure 2: Condition on the ratio $M_{\bullet}/M_{\star}$ for quasar feedback to suppress the average star formation efficiency. Active galaxies that reside in the green area, with a ratio $M_{\bullet}/M_{\star}$ above the threshold (indicated with a black line), experience quasar activity that increases the gas temperature above the virial value. Colored symbols indicate overmassive systems discovered by JWST at $z>4$, as the legend indicates. Gray symbols indicate local galaxies on the $M_{\bullet}-M_{\star}$ relation from Reines & Volonteri (2015), whose ratio $M_{\bullet}/M_{\star}$ is shown as a dashed line.
@@ 3.2 The Redshift Evolution of the $M_{\bullet}-M_{\star}$ Relation
$\frac{M_{\bullet}}{M_{\star}}\propto\xi(z)^{5/6}(1+z)^{5/2}$ (7)
@@ 4 Discussion and Conclusions
Significant uncertainties affect the determination of the stellar mass, derived from SED fitting to galaxy templates, and the SMBH mass, derived from single-epoch virial estimators, based, for example, on the width of the $\rm H\alpha$ line of the broad line region (see, e.g., Greene & Ho 2005). These methods are calibrated in the local Universe ($z\ll 1$) and have yet to be thoroughly tested at higher redshift (see, e.g., the discussion in Maiolino et al. 2023a). Notwithstanding these uncertainties, to retrieve the local scaling relations, the black hole masses (stellar masses) of these high-$z$ overmassive systems would need to be overestimated (underestimated) by a factor $\sim 60$ (Pacucci et al., 2023).
@@
Our model suggests that early SMBHs, primarily if formed as heavy seeds of initial mass $\sim 10^{4}-10^{5}\,{\rm M_{\odot}}$ (see, e.g., Ferrara et al. 2014), affect the evolution of the entire host. Eventually, the activity duty cycle of the quasar drops significantly below unity, and efficient star formation can resume. Once the galaxy grows via mergers, more stars and cool gas are added. Eventually, stars catch up with the SMBH mass, self-regulation of star formation occurs (see, e.g., Wyithe & Loeb 2003), and the system reaches the local $M_{\bullet}-M_{\star}$ relation.
Understanding the high-redshift evolution of the scaling relations is fundamental for two reasons. First, it informs us about the physical processes that regulate the growth of the black hole and stellar component (see, e.g., Vogelsberger et al. 2014; Schaye et al. 2015; Weinberger et al. 2017; Nelson et al. 2018; Terrazas et al. 2020; Piotrowska et al. 2022; Bluck et al. 2023). Second, it may inform us of the seeding mechanism that formed the central black hole in the first place. In fact, several studies have shown that a high ratio $M_{\bullet}/M_{\star}$ may be indicative of the formation of a heavy seed (see, e.g., Agarwal et al. 2013; Natarajan et al. 2017; Visbal & Haiman 2018; Scoggins et al. 2023; Natarajan et al. 2023) at $z>20$. The study of the properties of central SMBHs and their hosts at high-$z$, as well as the detection of extremely massive, and rare SMBHs at $z>10$, will determine if heavy seed formation channels were active in the high-$z$ Universe (Pacucci & Loeb, 2022).

### 03 Mar 2024 v2

@@ 1 Introduction
Other works have studied the redshift evolution of scaling relations using numerical simulations (e.g., Robertson et al. 2006; Di Matteo et al. 2008; Sijacki et al. 2015), semi-analytic models (e.g., Malbon et al. 2007; Kisaka & Kojima 2010), observations (e.g., Peng et al. 2006; Decarli et al. 2010; Merloni et al. 2010; Trakhtenbrot & Netzer 2010; Bennert et al. 2011), or combinations of those (e.g., Booth & Schaye 2011), especially at $z\lesssim 2$. More recently, Caplar et al. (2018) developed a phenomenological model, based on observations, and found that the ratio $M_{\bullet}/M_{\star}$ scales as $\propto(1+z)^{1.5}$ at $z<2$, in agreement with Wyithe & Loeb (2003). Furthermore, using simple but effective arguments on the density and dust content of the high-$z$ population of galaxies detected by JWST, very recently Inayoshi & Ichikawa (2024) proposed a high-$z$ $M_{\bullet}-M_{\star}$ that is in complete agreement with ours.
@@ 2 Model Principles
The basis of our model is that the black hole mass is the primary parameter that controls high-$z$ star formation quenching. This premise is supported by recent analyses of JWST/CEERS data with cosmological simulations (Illustris TNG and EAGLE), showing that high-$z$ galaxy quenching is primarily regulated by the mass of the SMBH (Piotrowska et al., 2022; Bluck et al., 2024). Previous cosmological simulations already showed that star formation quenching should scale with energy input from the central SMBH over the entire lifetime of the galaxy, which is proportional to the black hole mass (Terrazas et al., 2020; Bluck et al., 2024). Active SMBHs in $z>4$ galaxies discovered by JWST are effective at quenching star formation for at least two reasons.
First, high-$z$ galaxies are physically small; the ionized bubble generated by active SMBHs likely extends to the entire galaxy. Typical physical sizes of galaxies detected by JWST at $z=7-9$ are characterized by effective radii of $80\,\mathrm{pc}<r_{e}<300\,\mathrm{pc}$, with a mean value of $r_{e}\sim 150$ pc (Baggen et al., 2023). Recently, Baldwin et al. (2024) estimated the size of GN-z11 as $150\pm 25$ pc. This typical physical size has to be compared with the radius of the ionization bubble created by a SMBH accreting at its Eddington rate. For a $\sim 10^{7}\,{\rm M_{\odot}}$ SMBH, as typically found in these overmassive systems (Pacucci et al., 2023), this radius can extend as much as $\sim 700$ kpc (see, e.g., Cen & Haiman 2000; Madau & Rees 2000; White et al. 2003). Regions of high-density gas presumably present in the high-$z$ galaxies could effectively shield the radiation from the SMBH and still allow localized star formation; this would, however, be ineffective in generating large-scale star formation. Hence, SMBHs were effective in quenching star formation because they had a global impact on the entire host. Note that previous studies (e.g., Chen et al. 2020) already highlighted the importance of the radius of star-forming galaxies in determining the growth of their SMBHs.
@@
Second, in the high-$z$ Universe, the growth time of black holes is comparable to the age of the Universe. Hence, central SMBHs at a given redshift $z$ are active for a time comparable to the Hubble time $t_{age}(z)$, if $z>4$. Figure 1 shows, for the high-$z$ overmassive systems detected thus far by JWST, a comparison between the age of the Universe (at detection) and growth time. This latter time is calculated assuming a continuous growth from a seeding redshift of $z=25$ (see, e.g., Barkana & Loeb 2001) at the Eddington rate, assuming a light seed of $100\,{\rm M_{\odot}}$ (left panel) or a heavy seed of $10^{4}\,{\rm M_{\odot}}$ (right panel). At higher redshift, the age of the Universe is comparable to, or even shorter than, the growth time (for those two specific seeding scenarios); hence, the “activity duty cycle” for those specific SMBHs has to be close to unity. An example to clarify our point. At the median redshift $z=5$ (Pacucci et al., 2023), the age of the Universe is $\sim 1.1$ Gyr. To reach a typical black hole mass of $10^{8}\,{\rm M_{\odot}}$, accreting at the Eddington rate, a light seed of $10^{2}\,{\rm M_{\odot}}$ would take $63\%$ of the age of the Universe (i.e., $692$ Myr); a heavy seed of $10^{5}\,{\rm M_{\odot}}$ would take $31\%$ of the age of the Universe (i.e., $346$ Myr). These black holes had to be “on” for a fraction of the age of the Universe close to unity, independent of the seeding mechanism. The seed mass is an important parameter but affects the growth time only logarithmically. In the example above, three orders of magnitude change in the seed mass affects the growth time only by a factor $\sim 2$. Such factors can be crucial in the very high-$z$ Universe, and to form extremely massive SMBHs. However, in the present study, we are describing SMBHs at typical redshifts of $\sim 5$. Hence, we argue that our model principles do not depend on the particular flavor of the black hole seed chosen.
Because their duty cycle is close to unity, these SMBHs are constantly injecting energy into the primeval galaxy and heating the cold gas necessary to produce stars. Once the age of the Universe is $\gtrsim 1$ Gyr, the SMBH is active only for a fraction of the Hubble time and stars can then form from cold molecular gas, which becomes widely available in the galaxy. This hypothesis is further confirmed by a recent analysis of $124.5<z<$ 12 galaxies in the JWST/CEERS survey (Cole et al., 2023), showing a higher variability of star formation activity at high redshift. Stars form primarily in short periods of starburst activity, with star-forming duty cycles of only $20\%$ at $z\sim 9$, and $40\%$ at $z\sim 5$. The study also suggests a smoother star formation activity at $z<4.5$, when the age of the Universe is $>1\,\rm Gyr$ and the quasar duty cycle drops significantly below unity.
@@ 3.1 A Condition for Star Formation Quenching
We assume that star formation is quenched once the SMBH injects sufficient energy into the system to raise the temperature above the virial one. Quasar feedback in the form of heating, not mechanical outflows, is thus responsible for quenching the average star formation efficiency; a multitude of studies have investigated the interplay between quasar feedback (both thermal and mechanical) and the formation of stars, both in the local and the high-$z$ Universe (see, e.g., Silk & Rees 1998; King 2003; Hickox et al. 2009; Cattaneo et al. 2009; Inayoshi & Haiman 2014; Weinberger et al. 2017). Recently, Gelli et al. (2023) used a similar formalism to argue that supernova feedback fails to quench star formation in high-$z$ galaxies. This finding supports our model principles detailed in Sec. 2.
@@
We adopt $\,{f_{\rm Edd}}=1$ for our calculations because overmassive systems discovered by JWST are estimated to be accreting at rates $50.1<\,{f_{\rm Edd}}<$ 5 (Harikane et al., 2023; Maiolino et al., 2023a). In particular, the Eddington ratio distribution of sources described in those two studies is skewed towards higher values, and well described by the statistics $\,{f_{\rm Edd}}=0.9^{+1.4}_{-0.3}$, where $0.9$ is the mean, and the upper and lower bounds are derived from the interquartile ranges. In this simple model, we implicitly assume that the fraction of the energy emitted by the quasar that is retained by the gas, $F_{q}$, is constant and does not depend on the stellar mass (see, e.g., Ferrarese 2002; Wyithe & Loeb 2003; Begelman 2004). The role of $F_{q}$ is similar to, and degenerate with, the Eddington ratio. In terms of heating, a higher fraction of energy retained by the gas would play the same role as a higher Eddington ratio. Current data do not warrant a more complex interdependence between $F_{q}$ and $M_{\star}$.
Furthermore, we use the median cooling function for the gas metallicity range $0.30.1<Z/Z_{\odot}<$ 0.3 and the mean gas number density ($\sim 0.5\,\mathrm{cm^{-3}}$) calculated for simulated galaxies in the redshift range $105<z<$ 10 by Robinson et al. (2022). The metallicity range used is justified by a recent study with JWST of $z\sim 6$ galaxies with masses $\sim 10^{10}\,{\rm M_{\odot}}$, showing typical values $12+\log(\mathrm{O/H})\sim 8.2$, which is $\approx 25\%$ of the solar value (Nakajima et al., 2023). Finally, we assume the values of the cosmological parameters from Planck Collaboration et al. (2020) and the halo mass to stellar mass ratio from Behroozi et al. (2019).
@@
Figure 2: Condition on the ratio $M_{\bullet}/M_{\star}$ for quasar feedback to suppress the average star formation efficiency. Active galaxies that reside in the green area, with a ratio $M_{\bullet}/M_{\star}$ above the threshold (indicated with a black line), experience quasar activity that increases the gas temperature above the virial value. Colored symbols indicate overmassive systems discovered by JWST at $z>4$, as the legend indicates. Gray symbols indicate local galaxies on the $M_{\bullet}-M_{\star}$ relation from Reines & Volonteri (2015), whose ratio $M_{\bullet}/M_{\star}$ is shown as a dashed line. The dotted lines indicate how the threshold ratio is affected by the range of variability of Eddington ratios in the high-$z$ sample considered; this is estimated with interquartile ranges, to properly describe skewness.
@@ 3.2 The Redshift Evolution of the $M_{\bullet}-M_{\star}$ Relation
$\frac{M_{\bullet}}{M_{\star}}\propto\xi(z)^{5/6}(1+z)^{5/2}\,.$ (7)
@@
This redshift evolution of the median value of the $M_{\bullet}-M_{\star}$ relation is based on the assumptions that $M_{\bullet}\propto v_{c}^{5}$ and $M_{\star}\propto v_{c}^{0}$; Section 2 describes the foundations upon which these assumptions are built. Of course, a shallower relation between black hole mass and circular velocity, and/or a positive, non-zero dependence between the stellar mass and the circular velocity, would lead to a milder redshift evolution, similar to what was found by other studies. For example, Caplar et al. (2018) derived phenomenologically that $M_{\bullet}/M_{\star}\propto(1+z)^{1.5}$. Previous studies (e.g., Decarli et al. 2010; Bennert et al. 2011) also found milder redshift evolutions of the ratio between black hole mass and host stellar masses; for example, the latter study found $M_{\bullet}/M_{\star}\propto(1+z)^{1.15}$. Further data at high redshift will clarify the redshift evolution of this fundamental relation.
3.2.1 Note on the Scatter Around the Relation
Our model correctly reproduces the redshift evolution of the normalization of the $M_{\bullet}-M_{\star}$ relation. Different evolution histories of the single galaxies cause the scatter around the relation; it may be due to second-order effects, such as the specifics of the accretion and merger histories of the single systems, as well as the fact that quasar feedback is likely anisotropic, and the gas distribution non-homogeneous. The prediction of these second-order effects is beyond the scope of this model and requires detailed numerical simulations of the single high-$z$ galactic systems.
Pacucci et al. (2023) performed an accurate statistical analysis of the relation between $M_{\bullet}$ and $M_{\star}$ for galaxies detected by JWST in the redshift range $74<z<$ 7. In particular, we adopted the likelihood function defined in Hogg et al. (2010), which is appropriate for data characterized by a relation that is “near-linear but not narrow, so there is an intrinsic width or scatter in the true relationship”. To account for this intrinsic scatter, one of the three parameters used to describe the inferred relation is $\nu$, which is the orthogonal variance of the best-fit intrinsic scatter. The standard orthogonal scatter in this formalism is $\sqrt{\nu}\sec(\theta)$, where $\theta$ is the angle between the inferred line and the horizontal axis. In the original paper, with 21 data points, the standard scatter is estimated as $0.69$ dex (Pacucci et al., 2023).
A new run of the algorithm described in Pacucci et al. (2023), to include all the $35$ galaxies used in the present study (an increase of $67\%$ in data points), led to the following results. First, the slope and intercept of the new $M_{\bullet}-M_{\star}$ relation are consistent with what found in Pacucci et al. (2023): $b=-2.54\pm 0.75$ (instead of $b=-2.43\pm 0.83$) and $m=1.12\pm 0.08$ (instead of $m=1.06\pm 0.09$). The standard orthogonal scatter is $0.53$ dex instead of $0.69$ dex. The scatter decreases because more data points are added in the higher-mass regions of the plot. Despite adding $67\%$ more data points compared to Pacucci et al. (2023), the intrinsic scatter decreases only by $23\%$. This example suggests the presence of an intrinsic width or scatter in the true relationship, which is large and due to the single evolutionary histories of the galaxies. Our model aims to characterize the average evolution of the population and cannot describe this scatter.
@@ 4 Discussion and Conclusions
Significant uncertainties affect the determination of the stellar mass, derived from SED fitting to galaxy templates, and the SMBH mass, derived from single-epoch virial estimators, based, for example, on the width of the $\rm H\alpha$ line of the broad line region (see, e.g., Greene & Ho 2005). These methods are calibrated in the local Universe ($z\ll 1$) and have yet to be thoroughly tested at higher redshift (see, e.g., the discussion in Maiolino et al. 2023a). Notwithstanding these uncertainties, to retrieve the local scaling relations, the black hole masses (stellar masses) of these high-$z$ overmassive systems would need to be overestimated (underestimated) by a factor $\sim 60$ (Pacucci et al., 2023). Aside from uncertainties on the data side, our model relies on simple, but not simplistic, assumptions that further data and simulations may prove to be incorrect, e.g., specific scaling relations for $M_{\bullet}$ and $M_{\star}$ with the circular velocity $v_{c}$, a specific relation for $v_{c}=v_{c}(M_{h},z)$, a limited (but realistic) range of Eddington ratios and a fixed fraction $F_{q}$ of quasar energy trapped by the gas within the galaxy.
@@
Our model suggests that early SMBHs, primarily if formed as heavy seeds of initial mass $\sim 10^{4}-10^{5}\,{\rm M_{\odot}}$ (see, e.g., Ferrara et al. 2014), affect the evolution of the entire host. Eventually, the activity duty cycle of the quasar drops significantly below unity, and efficient star formation can resume. Once the galaxy grows via mergers, more stars and cool gas are added. Eventually, stars catch up with the SMBH mass, self-regulation of star formation occurs (see, e.g., Wyithe & Loeb 2003), and the system reaches the local $M_{\bullet}-M_{\star}$ relation. Note that the redshift evolution of the $M_{\bullet}-M_{\star}$ relation does not prevent the existence of outliers in the mass distribution, both at low and at high redshifts. Our model describes the redshift evolution of the median relation.
Understanding the high-redshift evolution of the scaling relations is fundamental for two reasons. First, it informs us about the physical processes that regulate the growth of the black hole and stellar component (see, e.g., Vogelsberger et al. 2014; Schaye et al. 2015; Weinberger et al. 2017; Nelson et al. 2018; Terrazas et al. 2020; Piotrowska et al. 2022; Bluck et al. 2024). Second, it may inform us of the seeding mechanism that formed the central black hole in the first place. In fact, several studies have shown that a high ratio $M_{\bullet}/M_{\star}$ may be indicative of the formation of a heavy seed (see, e.g., Agarwal et al. 2013; Natarajan et al. 2017; Visbal & Haiman 2018; Scoggins et al. 2023; Natarajan et al. 2024) at $z>20$. The study of the properties of central SMBHs and their hosts at high-$z$, as well as the detection of extremely massive, and rare SMBHs at $z>10$, will determine if heavy seed formation channels were active in the high-$z$ Universe (Pacucci & Loeb, 2022).